# Runtime Analysis of the SMS-EMOA for Many-Objective Optimization

### v2

@@ Abstract
This paper conducts the first rigorous runtime analysis of the SMS-EMOA for many-objective optimization. To this aim, we first propose a many-objective counterpart, the $m$-objective $m$OJZJ problem, of the bi-objective OJZJ benchmark, which is the first many-objective multimodal benchmark used in a mathematical runtime analysis. We prove that SMS-EMOA computes the full Pareto front of this benchmark in an expected number of $O(M^{2}n^{k})$ iterations, where $n$ denotes the problem size (length of the bit-string representation), $k$ the gap size (a difficulty parameter of the problem), and $M=(2n/m-2k+3)^{m/2}$ the size of the Pareto front. This result together with the existing negative result on the original NSGA-II shows that in principle, the general approach of the NSGA-II is suitable for many-objective optimization, but the crowding distance as tie-breaker has deficiencies.
@@
We obtain three additional insights on the SMS-EMOA. Different from a recent result for the bi-objective OJZJ benchmark, the stochastic population update often does not help for $m$OJZJ. It results in a $1/\Theta(\min\{Mk^{1/2}/2^{k/2},1\})$ speed-up, which is $\Theta(1)$ for large $m$ such as $m>k$. On the positive side, we prove that heavy-tailed mutation still results in a speed-up of order $k^{0.5+k-\beta}$. Finally, we conduct the first runtime analyses of the SMS-EMOA on the bi-objective OneMinMax and LOTZ benchmarks and show that it has a performance comparable to the GSEMO and the NSGA-II.
Introduction
@@
The NSGA-II is the most widely-applied multiobjective evolutionary algorithm (MOEA). Non-dominated sorting and crowding distance are its two major features differentiating it from basic MOEAs such as the GSEMO or the $(\mu+1)$ SIBEA. Zheng, Liu, and Doerr (2022) conducted the first runtime analysis of the NSGA-II (see (Zheng and Doerr 2023a) for the journal version). This work quickly inspired many interesting follow-up results in bi-objective optimization (Zheng and Doerr 2022; Bian and Qian 2022; Doerr and Qu 2023a, b, c; Dang et al. 2023a, b; Cerf et al. 2023). In contrast to these positive results for two objectives, Zheng and Doerr (2023b) proved that for $m\geq 3$ objectives the NSGA-II needs at least exponential time (in expectation and with high probability) to cover the full Pareto front of the $m$-objective OneMinMax benchmark, a simple many-objective version of the basic OneMax problem where all search points are Pareto optimal. They claimed that the main reason for this low efficiency is the independent computation of the crowding distance in each objective.
@@
A very recent work showed that the NSGA-III, a successor algorithm of the NSGA-II aimed at better coping with many objectives, can efficiently solve the $3$-objective OneMinMax problem (Wietheger and Doerr 2023). Since apparently practitioners much prefer the NSGA-II (more than 17,000 citations on Google scholar only in the last five years) over the NSGA-III (5,035 citations since its publication in 2013), it remains an interesting question whether there are variants of the NSGA-II which better cope with many objectives.
@@
With the SMS-EMOA, an interesting variant of the NSGA-II was proposed by Beume, Naujoks, and Emmerich (2007). This algorithm is a steady-state variant of the NSGA-II (that is, in each iteration only a single offspring is generated and possibly integrated into the population) that further replaces the crowding distance as secondary selection criterion with the classic hypervolume contribution. Many empirical works (see the almost 2,000 papers citing (Beume, Naujoks, and Emmerich 2007)) confirmed the good performance of the SMS-EMOA for many-objective optimization. The first mathematical runtime analysis of the SMS-EMOA was conducted very recently by Bian et al. (2023), who proved that its expected runtime on the bi-objective OJZJ problem is $O(n^{k+1})$. They also proposed a stochastic population update mechanism and proved that it has the often superior runtime of $O(n^{k+1}\min\{1,n/2^{k/4}\})$. Zheng et al. (2024) proved that the SMS-EMOA has an expected runtime of $O(n^{4})$ on the bi-objective DLTB problem.
@@
Our Contributions: This paper conducts the first mathematical runtime analysis of the SMS-EMOA for more than two objectives. We first define the $m$OJZJ benchmark, an $m$-objective counterpart of the bi-objective OJZJ problem (Zheng and Doerr 2023d), which is the problem analyzed in the first runtime analysis for the SMS-EMOA (Bian et al. 2023). We note that the $m$OJZJ problem is the first multimodal many-objective benchmark proposed for a theoretical analysis, to the best of our knowledge. We prove that the SMS-EMOA covers the full Pareto front of this benchmark in an expected number of $O(M^{2}n^{k})$ iterations, where $n$ is the problems size, $k$ the gap size (a difficulty parameter of the problem), $m$ the number of objectives, and $M=(2n/m-2k+3)^{m/2}$ the size of the Pareto front.
@@
We recall that the original NSGA-II needs at least exponential time to optimize the $m$OneMinMax problem, which is a special case of $m$OJZJ with gap size $k=1$. Since the SMS-EMOA employs non-dominated sorting, but replaces the crowding distance in the original NSGA-II by the hypervolume contribution, our result in a similar fashion as the analysis of the NSGA-III in (Wietheger and Doerr 2023) suggests that the general approach of the NSGA-II is suitable for many-objective optimization and that it is only the crowding distance as tie-breaker which is not appropriate for more than two objectives.
@@
We then analyze whether the better performance of the SMS-EMOA on the bi-objective OJZJ problem achieved via a new stochastic population update (Bian et al. 2023) extends to the $m$-objective $m$OJZJ problem. Unfortunately, we shall observe that only a speed-up of order $1/\Theta(\min\{Mk^{1/2}/2^{k/2},1\})$ is obtained, which is $\Theta(1)$ when $m$ is large, e.g., $m>k$.
@@
On the positive side, we show that the advantage of heavy-tailed mutation is preserved. We analyze the SMS-EMOA with heavy-tailed mutation on the $m$OJZJ benchmark and prove that a speed-up of order $k^{0.5+k-\beta}$ is achieved. This is the same speed-up as observed for single-objective and bi-objective Jump problems (Doerr et al. 2017; Zheng and Doerr 2023d). We note that this is the first theoretical work to support the usefulness of heavy-tailed mutation in many-objective optimization.
@@
Finally, since so far the performance of the SMS-EMOA was only analyzed on the bi-objective OJZJ problem (Bian et al. 2023), we conduct mathematical runtime analyses of the SMS-EMOA also on the two most prominent bi-objective benchmarks OneMinMax and LOTZ. We prove that the SMS-EMOA finds the Pareto fronts of these benchmarks in an expected number of iterations of at most $2e(n+1)n(\ln n+1)$ for OneMinMax and at most $2en^{2}(n+1)$ for LOTZ. These are the same asymptotic runtimes (in terms of fitness evaluations) as known for the GSEMO and the NSGA-II.
Preliminaries
@@
In many-objective optimization, one tries to find good solutions for a problem containing several, but at least three objectives. This work considers pseudo-Boolean maximization, hence our problem is described by a function $f=(f_{1},\dots,f_{m}):\{0,1\}^{n}\rightarrow\mathbb{R}^{m}$. Here $m>2$ is the number of objectives and $n\in\mathbb{N}$ is called the problem size. As in bi-objective optimization (that is, the case $m=2$), usually not all solutions are comparable. We say that a solution $x\in\{0,1\}^{n}$ dominates a solution $y$, denoted as $x\succ y$, if $f_{i}(x)\geq f_{i}(y)$ for all $i\in\{1,\dots,m\}$, and at least one of these inequalities is strict. For a given many-objective function $f$, we say that $x\in\{0,1\}^{n}$ is a Pareto optimum if and only if there is no $y\in\{0,1\}^{n}$ that dominates $x$ w.r.t. $f$. The set of all Pareto optima is called Pareto set, and the set of all function values of the Pareto optima is called Pareto front. The goal for an MOEA is to find the full Pareto front as much as possible, that is, to compute a not too large set $S$ of solutions such that $f(S)$ equals or approximates well the Pareto front. As common in the mathematical runtime analysis of MOEAs, we call the runtime of an algorithm the number of function evaluations until its population $P$ covers the full Pareto front, that is, $f(P)$ contains the Pareto front. We refer to (Neumann and Witt 2010; Auger and Doerr 2011; Jansen 2013; Zhou, Yu, and Qian 2019; Doerr and Neumann 2020) for general introductions to the mathematical runtime analysis of evolutionary algorithms, and to (Brockhoff 2011) for a discussion of runtime analyses of MOEAs.
@@
Mathematical Notations. For $a,b\in\mathbb{N}$ with $a\leq b$, we denote the set of integers in the interval $[a,b]$ by $[a..b]=\{a,a+1,\dots,b\}$. For a point $x\in\{0,1\}^{n}$, let $x_{[a..b]}:=(x_{a},x_{a+1},\dots,x_{b})$ and $|x|_{1}$ be the number of ones in $x$.
The $m$-Objective Jump Function
@@
As mentioned earlier, the NSGA-II was proven to have enormous difficulties in optimizing many-objective problems (Zheng and Doerr 2023b). In that paper, the $m$-objective counterpart $m$OneMinMax of the bi-objective OneMinMax benchmark was proposed and analyzed. Since the first runtime work of the SMS-EMOA so far (Bian et al. 2023) analyzed the bi-objective jump problem OJZJ, whose special case with gap size $k=1$ is (essentially) the OneMinMax problem, we shall propose and work with an $m$-objective version of OJZJ as well. Again, its special case $k=1$ will be (essentially) equal to the $m$OneMinMax problem. With this, our results are comparable both the ones in (Zheng and Doerr 2023b) and (Bian et al. 2023).
$m$OJZJ
@@
We first recall the definition of the $m$OneMinMax problem. For the ease of presentation, we only consider even numbers $m$ of objectives here. In the $m$OneMinMax problem, the bit string (of length $n$) is divided into $m/2$ blocks of equal length $2n/m$. On each of these, a bi-objective OneMinMax problem is defined. We note that this general block construction goes back to the seminal paper of (Laumanns, Thiele, and Zitzler 2004).
Definition 1 (Zheng and Doerr (2023b)).
@@
Let $m$ be the number of objectives and be even, and the problem size $n$ be a multiple of $m/2$. Let $n^{\prime}=2n/m\in\mathbb{N}$. For any $x=(x_{1},\dots,x_{n})$, the $m$-objective function mOneMinMax $f:\{0,1\}^{n}\rightarrow\mathbb{R}^{m}$ is defined by
@@
$f_{i}(x)$ =$=\begin{cases}\textsc{OneMax}(\bar{x}_{[\frac{i-1}{2}n^{\prime}+1..\frac{i+1}{2}n^{\prime}]}),&\text{if $i$ is odd}\\ \textsc{OneMax}({x}_{[\frac{i-2}{2}n^{\prime}+1..\frac{i}{2}n^{\prime}]}),&\text{else,}\end{cases}$
@@
In order to include $m$OneMinMax as a special case of $m$OJZJ, we define the $m$-objective OJZJ in a similar manner, that is, we divide the $n$ bit positions into $m/2$ blocks and define a OJZJ problem in each block.
@@
$f_{i}(x)$ =$=\begin{cases}\textsc{Jump}_{n^{\prime},k}(x_{[\frac{i-1}{2}n^{\prime}+1..\frac{i+1}{2}n^{\prime}]}),&\text{if $i$ is odd}\\ \textsc{Jump}_{n^{\prime},k}(\bar{x}_{[\frac{i-2}{2}n^{\prime}+1..\frac{i}{2}n^{\prime}]}),&\text{else,}\end{cases}$
@@
for any $y\in\{0,1\}^{n^{\prime}}$.
We note that the function $\textsc{Jump}_{n^{\prime},k}$ used in the definition above is the famous Jump benchmark. It was first defined by Droste, Jansen, and Wegener (2002) and has quickly become the most employed multimodal benchmark in the theory of randomized search heuristics, leading to many fundamental results on how these algorithms cope with local optima (Jansen and Wegener 2002; Dang et al. 2018; Doerr 2021; Benbaki, Benomar, and Doerr 2021; Hevia Fajardo and Sudholt 2022; Rajabi and Witt 2022; Witt 2023; Doerr et al. 2024).
[…]
Characteristics
We now give more details on this $m$OJZJ function. Let $B_{i}:=[(i-1)n^{\prime}+1..in^{\prime}]$, $i\in[1..m/2]$, be the $i$-th block in the partition of $n$ bit positions. From Definition 2, we know that the bit values in the block $B_{i}$ only influence the objectives $f_{2i-1}$ and $f_{2i}$. Figure 1 plots the objective values of $f_{2i-1}$ and $f_{2i}$ relative to the number of ones in this block. Obviously, $m$OJZJ is multimodal with respect to the definition of multimodality of multiobjective problems in (Zheng and Doerr 2023d).

### v3

@@ Abstract
This paper conducts the first rigorous runtime analysis of the SMS-EMOA for many-objective optimization. To this aim, we first propose a many-objective counterpart, the $m$-objective $m$OJZJ problem, of the bi-objective OJZJ benchmark, which is the first many-objective multimodal benchmark used in a mathematical runtime analysis. We prove that SMS-EMOA computes the full Pareto front of this benchmark in an expected number of $O(\mu Mn^{k})$ iterations, where $n$ denotes the problem size (length of the bit-string representation), $k$ the gap size (a difficulty parameter of the problem), $M=(2n/m-2k+3)^{m/2}$ the size of the Pareto front, and $\mu$ the population size (at least the same size as the largest incomparable set). This result together with the existing negative result for the original NSGA-II shows that, in principle, the general approach of the NSGA-II is suitable for many-objective optimization, but the crowding distance as tie-breaker has deficiencies.
@@
We obtain three additional insights on the SMS-EMOA. Different from a recent result for the bi-objective OJZJ benchmark, a recently proposed stochastic population update often does not help for $m$OJZJ. It at most results in a speed-up by a factor of order $2^{k}/\mu$, which is $\Theta(1)$ for large $m$, such as $m>k$. On the positive side, we prove that heavy-tailed mutation irrespective of the number $m$ of objectives results in a speed-up of order $k^{0.5+k-\beta}/e^{k}$, the same advantage as previously shown for the bi-objective case. Finally, we conduct the first runtime analyses of the SMS-EMOA on the classic bi-objective OneMinMax and LOTZ benchmarks and show that the SMS-EMOA has a performance comparable to the GSEMO and the NSGA-II.
Our main technical insight, a general condition ensuring that the SMS-EMOA does not lose Pareto-optimal objective values, promises to be useful also in other runtime analyses of this algorithm.
1 Introduction
@@
The NSGA-II is the most widely-applied multiobjective evolutionary algorithm (MOEA). Non-dominated sorting and crowding distance are its two main building blocks differentiating it from basic MOEAs such as the GSEMO or the $(\mu+1)$ SIBEA. Zheng, Liu, and Doerr [ZLD22] conducted the first runtime analysis of the NSGA-II (see [ZD23a] for the journal version). This work quickly inspired many interesting follow-up results in bi-objective optimization [ZD22, BQ22, DQ23a, DQ23b, DQ23c, DOSS23a, DOSS23b, CDH+23]. In contrast to these positive results for two objectives, Zheng and Doerr [ZD23b] proved that for $m\geq 3$ objectives the NSGA-II needs at least exponential time (in expectation and with high probability) to cover the full Pareto front of the $m$-objective OneMinMax benchmark, a simple many-objective version of the basic OneMax problem. They claimed that the main reason for this low efficiency is the independent computation of the crowding distance in each objective, which suggested that the NSGA-II would have similar difficulties on other many-objective problems.
@@
A very recent work showed that the NSGA-III, a successor algorithm of the NSGA-II aimed at better coping with many objectives, can efficiently solve the $3$-objective OneMinMax problem [WD23]. Since apparently practitioners much prefer the NSGA-II (more than 17,000 citations on Google scholar only in the last five years) over the NSGA-III (5,035 citations since its publication in 2013), it remains an interesting question whether there are other variants of the NSGA-II which also cope well with many objectives.
@@
With the SMS-EMOA, an interesting variant of the NSGA-II was proposed by Beume, Naujoks, and Emmerich [BNE07]. This algorithm is a steady-state variant of the NSGA-II (that is, in each iteration only a single offspring is generated and possibly integrated into the population) that further replaces the crowding distance as secondary selection criterion with the classic hypervolume contribution. Many empirical works (see the almost 2,000 papers citing [BNE07]) confirmed the good performance of the SMS-EMOA for many-objective optimization. The first mathematical runtime analysis of the SMS-EMOA was conducted very recently by Bian et al. [BZLQ23], who proved that its expected runtime on the bi-objective OJZJ problem is $O(n^{k+1})$. They also proposed a stochastic population update mechanism and proved that it has the often superior runtime of $O(n^{k+1}\min\{1,n/2^{k/4}\})$. Zheng et al. [ZLDD24] proved that the SMS-EMOA has an expected runtime of $O(n^{4})$ on the bi-objective DLTB problem.
@@
Our Contributions: This paper conducts the first mathematical runtime analysis of the SMS-EMOA for more than two objectives. We first define the $m$OJZJ benchmark, an $m$-objective counterpart of the bi-objective OJZJ problem [ZD23c], which is the problem analyzed in the first runtime analysis for the SMS-EMOA [BZLQ23]. We note that the $m$OJZJ problem is the first multimodal many-objective benchmark proposed for a theoretical analysis, to the best of our knowledge. We prove in Theorem 9 that the SMS-EMOA can compute the full Pareto front of this benchmark in an expected number of $O(\mu Mn^{k})$ iterations, where $n$ is the problems size, $k$ the gap size (a difficulty parameter of the problem that is usually small), $m$ the number of objectives, $M=(2n/m-2k+3)^{m/2}$ the size of the Pareto front, and where the population size $\mu$ has to be at least as large as the largest set of incomparable solutions (which is not larger than $(2n/m+1)^{m/2}$).
@@
We recall that the original NSGA-II needs at least exponential time to optimize the $m$OneMinMax problem, which is a special case of $m$OJZJ with gap size $k=1$. Since the SMS-EMOA employs non-dominated sorting, but replaces the crowding distance in the original NSGA-II by the hypervolume contribution, our result in a similar fashion as the analysis of the NSGA-III in [WD23] suggests that the general approach of the NSGA-II is suitable for many-objective optimization and that it is only the crowding distance as tie-breaker which is not appropriate for more than two objectives.
@@
We then analyze whether the better performance of the SMS-EMOA on the bi-objective OJZJ problem achieved via a new stochastic population update [BZLQ23] extends to the $m$-objective $m$OJZJ problem. Unfortunately, we shall observe that only a speed-up of order at most $2^{k}/\mu$ is obtained, which is $\Theta(1)$ when $m$ is large, e.g., $m>k$.
@@
On the positive side, we show that the advantage of heavy-tailed mutation is preserved. We analyze the SMS-EMOA with heavy-tailed mutation on the $m$OJZJ benchmark and prove that a speed-up of order $k^{0.5+k-\beta}/e^{k}$ is achieved. This is the same speed-up as observed for single-objective and bi-objective Jump problems [DLMN17, ZD23c]. We note that this is the first theoretical work to support the usefulness of heavy-tailed mutation in many-objective optimization.
@@
Finally, since so far the performance of the SMS-EMOA was only analyzed on the bi-objective OJZJ problem [BZLQ23] (and, in a work [ZLDD24] parallel to this, on the not overly prominent DLTB benchmark), we conduct mathematical runtime analyses of the SMS-EMOA also on the two most prominent bi-objective benchmarks OneMinMax and LOTZ. We prove that the SMS-EMOA finds the Pareto fronts of these benchmarks in an expected number of at most $2e(n+1)n(\ln n+1)$ iterations for OneMinMax and at most $2en^{2}(n+1)$ for LOTZ. These are the same asymptotic runtimes (in terms of fitness evaluations) as known for the GSEMO and the NSGA-II.
The technical key to all these results are two simple, but general results ensuring that the SMS-EMOA (different from the NSGA-II in many-objective optimization) cannot forget Pareto-optimal solution values when the population size is sufficiently large (Lemmas 4 and 20). We are convinced that these will find applications in future runtime analyses of the SMS-EMOA. In fact, Lemma 4 has already been used in the subsequent work [WD24].
This paper is organized as follows. Section 2 briefly introduces some basic concepts for the multi-objective optimization and the mathematical notations that will be used in this paper. The many-objective OJZJ benchmark is proposed in Section 3. Sections 4 to 6 analyze the runtime of the original SMS-EMOA, the SMS-EMOA with stochastic population update, and the one with heavy-tailed mutation on this many-objective benchmark. The runtime for two well-known bi-objective functions is conducted in Section 7. Section 8 concludes this work.
2 Preliminaries
In multi-objective optimization, one tries to find good solutions for a problem containing several objectives. This work considers pseudo-Boolean maximization, hence our problem is described by a function $f=(f_{1},\dots,f_{m}):\{0,1\}^{n}\rightarrow\mathbb{R}^{m}$. Here $m$ is the number of objectives and $n\in\mathbb{N}$ is called the problem size. We use the term bi-objective optimization when $m=2$, and many-objective optimization when $m\geq 3$. The literature is not uniform in from which number of objectives to speak of many objectives. We include already $m=3$ into the many-objective setting since the drastically different behavior of the NSGA-II detected in [ZD23b] starts already at $m=3$.
@@ 2 Preliminaries
Different from single-objective optimization, usually not all solutions are comparable in multi-objective problems. We say that a solution $x\in\{0,1\}^{n}$ dominates a solution $y$ (with regard to $f$), denoted as $x\succ y$, if $f_{i}(x)\geq f_{i}(y)$ for all $i\in\{1,\dots,m\}$, and at least one of these inequalities is strict. For a given multi-objective function $f$, we say that $x\in\{0,1\}^{n}$ is a Pareto optimum if and only if there is no $y\in\{0,1\}^{n}$ that dominates $x$ (with regard to $f$). The set of all Pareto optima is called Pareto set, and the set of all function values of Pareto optima is called Pareto front. The goal for an MOEA is to find the full Pareto front as good as possible, that is, to compute a not too large set $S$ of solutions such that $f(S)$ equals or approximates well the Pareto front. As common in the mathematical runtime analysis of MOEAs, we call the runtime of an algorithm the number of function evaluations until its population $P$ covers the full Pareto front, that is, $f(P)$ contains the Pareto front. We refer to [NW10, AD11, Jan13, ZYQ19, DN20] for general introductions to the mathematical runtime analysis of evolutionary algorithms, and to [Bro11] for a discussion of runtime analyses of MOEAs.
@@
Mathematical notation. For $a,b\in\mathbb{N}$ with $a\leq b$, we denote the set of integers in the interval $[a,b]$ by $[a..b]=\{a,a+1,\dots,b\}$. For a point $x\in\{0,1\}^{n}$, let $x_{[a..b]}:=(x_{a},x_{a+1},\dots,x_{b})$ and $|x|_{1}$ be the number of ones in $x$.
3 The $m$-Objective Jump Problem
@@
As mentioned earlier, the NSGA-II was proven to have enormous difficulties in optimizing many-objective problems [ZD23b]. In that paper, the $m$-objective counterpart $m$OneMinMax of the bi-objective OneMinMax benchmark [GL10] was proposed and analyzed. Since the first runtime work of the SMS-EMOA so far [BZLQ23] analyzed the bi-objective jump problem OJZJ, whose special case with gap size $k=1$ is (essentially) the OneMinMax problem, we shall propose and work with an $m$-objective version of OJZJ. Again, its special case $k=1$ will be (essentially) equal to the $m$OneMinMax problem. With this, our results are comparable both the ones in [ZD23b] and [BZLQ23].
3.1 $m$OJZJ
@@
We first recall the definition of the $m$OneMinMax problem. For the ease of presentation, we only consider even numbers $m$ of objectives here. In the $m$OneMinMax problem, the bit string (of length $n$) is divided into $m/2$ blocks of equal length $2n/m$. On each of these, a bi-objective OneMinMax problem is defined. We note that this general approach to construct many-objective problems from bi-objective ones goes back to the seminal paper of [LTZ04].
Definition 1 ([ZD23b]).
@@
Let $m$ be the even number of objectives. Let the problem size $n$ be a multiple of $m/2$. Let $n^{\prime}=2n/m\in\mathbb{N}$. For any $x=(x_{1},\dots,x_{n})$, the $m$-objective function mOneMinMax is the function $f:\{0,1\}^{n}\rightarrow\mathbb{R}^{m}$ defined by
@@
$f_{i}(x)$ =$=\begin{cases}\textsc{OneMax}(\bar{x}_{[\frac{i-1}{2}n^{\prime}+1..\frac{i+1}{2}n^{\prime}]}),&\text{if $i$ is odd,}\\ \textsc{OneMax}({x}_{[\frac{i-2}{2}n^{\prime}+1..\frac{i}{2}n^{\prime}]}),&\text{else,}\end{cases}$
@@
We define the $m$-objective OJZJ in a similar manner, that is, we divide the $n$ bit positions into $m/2$ blocks and define a OJZJ problem in each block.
@@
$f_{i}(x)$ =$=\begin{cases}\textsc{Jump}_{n^{\prime},k}(x_{[\frac{i-1}{2}n^{\prime}+1..\frac{i+1}{2}n^{\prime}]}),&\text{if $i$ is odd,}\\ \textsc{Jump}_{n^{\prime},k}(\bar{x}_{[\frac{i-2}{2}n^{\prime}+1..\frac{i}{2}n^{\prime}]}),&\text{else,}\end{cases}$
@@
$\textsc{Jump}_{n^{\prime},k}(y)$ =$=\begin{cases}k+\textsc{OneMax}(y),&\text{if $|y|_{1}\leq n^{\prime}-k$ or $y=1^{n^{\prime}}$}\\ n^{\prime}-\textsc{OneMax}(y),&\text{else}\end{cases}$
for any $y\in\{0,1\}^{n^{\prime}}$.
We note that the function $\textsc{Jump}_{n^{\prime},k}$ used in the definition above is the famous Jump benchmark. It was first defined by Droste, Jansen, and Wegener [DJW02] and has quickly become the most employed multimodal benchmark in the theory of randomized search heuristics, leading to many fundamental results on how these algorithms cope with local optima [JW02, DFK+18, Doe21, BBD21, HS22, RW22, Wit23, DEJK24].
@@
3.2 Characteristics
We now give more details on this $m$OJZJ function. Let $B_{i}:=[(i-1)n^{\prime}+1..in^{\prime}]$, $i\in[1..m/2]$, be the $i$-th block of the $n$ bit positions. From Definition 2, we know that the bit values in block $B_{i}$ only influence the objectives $f_{2i-1}$ and $f_{2i}$. Figure 1 plots the objective values of $f_{2i-1}$ and $f_{2i}$ relative to the number of ones in this block. Obviously, $m$OJZJ is multimodal with respect to the definition of multimodality of multiobjective problems in [ZD23c].